OpenStudy (anonymous):

How do I convert 7.59 into a mixed number

OpenStudy (anonymous):

don't just give me the answer, but show me how you got the answer.

OpenStudy (anonymous):

Do you know?

OpenStudy (mathmath333):

\(\Large\tt \begin{align} \color{black}{7.59 \\~\\ =7+0.59\\~\\ =7+\dfrac{59}{100}\\~\\ =7\cfrac{59}{100}\\~\\}\end{align}\)

OpenStudy (anonymous):

Thank you so much!!! :)